A city park is opening up a new swimming pool. you can pay a daily entrance fee of $3 or purchase a membership for the 12 week summer season for $82 and only pay $1 per day to swim. How many days would you have to swim to make the membership worth while?
Answer by rfer(16322)

3x=82+x
2x=82
x=41
--------------
3*41=82+41